# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

### v2

@@ Abstract
In this paper, we propose a coordinated routing strategy aimed at improving bus schedule adherence and enhancing travel efficiency for connected and automated vehicles (CAVs) operating within a mixed-traffic urban network. Our approach capitalizes on the existence of dedicated lanes for buses and CAVs, leveraging real-time traffic data to dynamically reroute CAVs in anticipation of congestion. By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, we enable the system to proactively adjust CAV routes when potential interference with bus operations is detected. This coordination mitigates delays affecting transit services and reduces travel time for CAVs. We evaluate the proposed strategy through simulation studies conducted in the SUMO. The results demonstrate significant improvements in both transit reliability and CAV operational performance across a range of traffic conditions.
@@ I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2]. While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [3, 4]. During this period, transportation networks will experience prolonged phases of mixed traffic, where CAVs and human-driven vehicles (HVs) coexist. Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [5, 6, 7]. Extensive research and experimental studies have shown that heterogeneous traffic environments involving CAVs can significantly reduce travel times [8, 9, 10]. Among various approaches, implementing suitable right-of-way management is pivotal for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transition.
@@
CAV dedicated lane (DL) has been proposed as an effective strategy to enhance the benefits that CAVs can bring to transportation systems [11]. By providing a separate driving environment, DLs enable CAVs to operate with shorter reaction times and smaller headways, thus increasing lane capacity by up to three times compared to general-purpose lanes (GPLs) [12]. In addition, DLs isolate CAVs from the mixed traffic network, offering a safer driving environment for both CAVs and HVs. This contributes to reducing vehicles’ lane changing and overtaking maneuvers, which are among the leading causes of collisions and delays in mixed traffic environments [13]. Despite the promising benefits of DLs, the high costs and limited feasibility of constructing new lanes in urban areas, due to substantial construction expenses and restricted urban space, often outweigh the advantages of CAV DLs [14]. In this context, converting an existing GPL into a CAV DL tends to be a more practical approach. However, it may lead to substantial delays for other types of vehicles using GPLs due to the reduced road capacity [15]. Recent research has shown that the overall system benefit from converting a GPL to a CAV DL may be lower than maintaining the original configuration [16].
@@
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [17]. While these lanes provide dedicated driving environments for transit operations, the reserved space cannot be taken by other non-transit vehicles, lowering the traffic efficiency [18]. Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability. In the existing literature, Luo et al. developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [19]. Chen et al. explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [20]. A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, was investigated in [12]. While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
@@
In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. The main contributions of this work are as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in SUMO based on realistic traffic network data in San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
@@ II Problem Statement
Figure 1: An example illustrating the road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
@@
Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As illustrated in Fig. 2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses. Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, hence improving road capacity while avoiding disturbance caused by other HVs on GPLs.
Figure 2: Illustration of the general-purpose lane and dedicated lane.
For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented as a sequence of edges as follows:
@@
For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ entering the road network, let $(o_{i},d_{i})$ denote its origin-destination (OD) pair, where $o_{i},d_{i}\!\in\!\mathcal{I}$. Based on real-time traffic conditions, CAV $i$ is assigned an optimal route from its location to the destination. The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted by
@@
where $v_{k},v_{k+1},\dots,v_{n}\!\in\!\mathcal{I}$ are road intersections in the pre-planned route of the CAV. Note that, the route of each CAV is dynamically updated upon its arrival at each intersection in its route. While CAV routes may traverse both intersections and bus stations, re-optimization of the route toward the destination is triggered only when approaching an intersection. HVs are treated as passive participants in the network, following certain routes during their trips.
@@ III Modeling of Traffic Flow
This section presents the modeling of road traffic flow in a dynamic transportation network. We start by introducing the traffic flow on joint DLs to capture the potential delays caused by CAVs. Then, the traffic flow on GPLs is modeled for effective rerouting of CAVs.
@@ III-A Traffic Flow on Joint Dedicated Lane
The travel time for a vehicle to traverse an edge depends on the real-time traffic flow along the edge. As the number of vehicles traveling on an edge increases, traffic flow becomes heavier and travel times tend to rise due to congestion. This nonlinear relationship is well captured by the classic Bureau of Public Roads (BPR) function [21]. Let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$. According to the BPR model, the actual travel time to traverse $e$ is denoted as
@@
where $f_{e}$ represents the real-time flow on edge $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion. Both $f_{e}$ and $c_{e}$ are measured in the number of vehicles per unit time. In this paper, we assume that all vehicles, regardless of type, travel at the same speed under free-flow conditions. Additionally, $\alpha$ and $\beta$ are congestion sensitivity parameters. Since CAVs react faster and drive closer with shorter headways, they usually smooth traffic flow compared to a fully human-driven lane [19]. This results in a lower value of $\alpha$. In contrast, buses are heavier and less flexible, causing greater congestion when traffic is dense [22]. Based on these considerations, we adopt the following parameter settings in this study
$\begin{cases}\alpha\!=\!0.1,\ \beta\!=\!3,&\text{for GPL},\\ \alpha\!=\!0.2,\ \beta\!=\!5,&\text{for joint DL}.\end{cases}$ (4)
Sensors deployed in the road network enable real-time monitoring and analysis of traffic flow patterns [23]. In the road network considered, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge. For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, the traffic flow monitor window is defined as
@@
where $t_{\text{sys}}$ denotes the current system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window where $\mathbb{N}_{+}$ denotes the set of positive integers. For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the transportation system at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its routing planning, which is represented as
@@ III-B Traffic Flow on General-Purpose Lane
Observing traffic flow on GPLs is essential to dynamically optimize CAV route planning based on real-time traffic data, thereby enhancing overall traffic efficiency. Similarly, the time window for traffic flow monitoring at the starting point of each edge on the GPLs is denoted as
@@
where $(v,v^{\prime})\!\in\!\mathcal{E}\!\setminus\!\cup_{b\in\mathcal{N}^{\text{cav}}}P_{b}$ denotes an edge located on the GPLs. Note that $\Delta{N}^{\text{hv}}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
Subsequently, according to the BPR model given in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right).$ (12)
Here, $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on edge $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow on the edge at time $t_{\text{sys}}$, computed according to (10). The congestion sensitivity parameters $\alpha$ and $\beta$ are determined by (4).
@@
The traffic flow monitoring windows are defined by $\Delta{T}$ in (5) for joint DLs and by $\Delta{\tilde{T}}$ in (9) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges. With a different emphasis, the sizes of the two monitoring windows are not necessarily identical.
@@ IV Dynamic Rerouting Approach
In this section, we propose a coordinated rerouting approach that dynamically optimizes route assignments for CAVs based on real-time traffic data. The objective is to ensure that buses adhere to their scheduled arrival times at each stop while enhancing the overall traffic efficiency of CAVs. The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and then determining the optimal alternative routes for those CAVs.
@@ IV-A Identification of CAVs for Rerouting
where $\tau_{b,k-1}^{0}$ denotes the free-flow travel time from intersection $v_{b,k-1}$ to intersection $v_{b,k}$.
@@
$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right),$ (15)
where $c_{b,k}$ is the effective capacity of edge $(v_{b,k},v_{b,k+1})$, defined as the maximum traffic flow that the road segment can accommodate under ideal conditions before congestion affects travel times. For edges on DLs, the parameters $\alpha$ and $\beta$ are chosen following the second case in (4).
Figure 3: Illustration of the rerouting scheme at each intersection in DLs.
To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$. The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected. We make use of Fig. 3 for better illustration.
For any system time $t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]$, the set of CAVs on edge $(v_{b,k-1},v_{b,k})$ that need to be rerouted, denoted as $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, is expressed as
@@
As defined (7), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of the intersection $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered. Namely,
$t_{\text{sys}}^{\prime}\in\big\{t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!{\tau_{b,k}^{0}}\big\}.$ (17)
Remark 2
In the developed rerouting approach, the traffic flow on edge $(v_{b,k},v_{b,k+1})$ is dynamically monitored, and the CAVs on the preceding edge, i.e., $(v_{b,k-1},v_{b,k})$, are controlled before bus $b$ entering $(v_{b,k},v_{b,k+1})$. This procedure enables effective congestion mitigation on each edge along the DLs. Thus, it is reasonable to employ the free-flow travel time in (14) to estimate the bus’s arrival time at $v_{b,k}$.

### v3

@@ Abstract
This paper proposes a coordinated routing approach that investigates the use of connected and automated vehicles (CAVs) in dedicated bus lanes. The aim is to improve bus schedule adherence while enhancing the travel efficiency of CAVs during the transitional phase of mixed traffic environments. Our approach utilizes real-time traffic data to dynamically reroute CAVs in anticipation of congestion. By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, the system adjusts CAV routes in advance to avoid potential interference with operating buses. This cooperation reduces CAV travel times and minimizes delays that impact transit services. The proposed strategy is validated using microscopic traffic simulations in SUMO. The results demonstrate significant improvements in both transit on-time performance and CAV travel efficiency across a range of traffic conditions.
@@ I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [12, 7, 8]. While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [13, 1]. Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [21, 11, 17]. Extensive research and experimental studies have shown that heterogeneous traffic environments involving CAVs can significantly reduce travel times [22, 9, 3]. Among various approaches, implementing suitable right-of-way management is pivotal for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transition.
@@
CAV dedicated lane (DL) has been proposed as an effective strategy to enhance the benefits that CAVs can bring to transportation systems [16]. By providing a separate driving environment, DLs enable CAVs to operate with shorter reaction times and smaller headways, thus increasing lane capacity by up to three times compared to general-purpose lanes (GPLs) [4]. However, in urban settings, the substantial construction costs and limited spatial availability often constrain the feasibility of implementing new CAV DLs [14]. In this context, converting an existing GPL into a CAV DL tends to be a more practical approach. However, it may lead to substantial delays for other types of vehicles using GPLs due to the reduced road capacity [10].
@@
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has emerged: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing human-driven vehicles (HVs), joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [18]. In the existing literature, Chen et al. proposed a space-time modeling framework demonstrating that the mixed use of DLs can enhance CAV travel efficiency and alleviate overall traffic congestion [5]. More recently, schedule-based joint optimization approaches have been presented to coordinate CAV speeds and bus timetables in Bus Rapid Transit corridors [20]. In addition, dynamic management strategies have been developed to allocate right-of-way between buses and CAVs through lane-level trajectory planning [19]. While these studies provide important insights into joint DL operations, they primarily focus on lane-level coordination or schedule-based optimization, and remain limited in dynamically regulating CAV routing at the network level, particularly in balancing congestion mitigation with reliable bus operations under time-varying demand.
@@
In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. In summary, the main contributions of this work are twofold. (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, which enables real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in the Simulation of Urban MObility (SUMO) environment based on realistic traffic network data of San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
@@ II Problem Statement
Figure 1: An example illustrating the road network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and DLs are shown by red edges.
@@
We consider three types of vehicles, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As given in Fig. 2, buses use only DLs and follow a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses.
For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented by a sequence of edges as follows:
@@
For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ entering the road network, let $(o_{i},d_{i})$ denote its origin-destination (OD) pair, where $o_{i},d_{i}\!\in\!\mathcal{I}$. Based on real-time traffic conditions, CAV $i$ is assigned an optimal route from its location to the destination. The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted as
@@
where $v_{k},v_{k+1},\dots,v_{n}\!\in\!\mathcal{I}$ are road intersections in the pre-planned route of the CAV. Note that the route of each CAV is dynamically updated upon its arrival at each intersection in its route. While CAV routes may traverse both intersections and bus stations, re-optimization of the route toward the destination is triggered only when approaching an intersection. HVs are treated as passive participants in the network, following certain routes during their trips.
Figure 2: A figure illustrating vehicle types, DLs, and GPLs.
@@ III Modeling of Traffic Flow
This section presents the modeling of the road traffic flow in a dynamic transportation network. We start by introducing the traffic flow on joint DLs to capture the potential delays caused by CAVs. Then, the traffic flow on GPLs is modeled for effective rerouting of CAVs.
@@ III-A Traffic Flow on Joint Dedicated Lane
The travel time for a vehicle to traverse an edge depends on the real-time traffic flow along the edge. Following the classic Bureau of Public Roads (BPR) function [6], let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$. The actual travel time on the edge is then modeled as
@@
where $f_{e}$ denotes the real-time flow on $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion. Both $f_{e}$ and $c_{e}$ are measured in the number of vehicles per unit time. We assume that all vehicles, regardless of their types, travel at the same speed under free-flow conditions. Additionally, $\alpha$ and $\beta$ are congestion sensitivity parameters.
Sensors deployed in the road network enable real-time monitoring and analysis of traffic flow patterns [2]. In the road network, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge. For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, we define the traffic flow monitor window
@@
where $t_{\text{sys}}$ denotes the system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window, with $\mathbb{N}_{+}$ denoting the set of positive integers. For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the road network at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its route planning, which is denoted as
@@ III-B Traffic Flow on General-Purpose Lane
Similarly, the time window for traffic flow monitoring at the starting point of each edge on the GPLs is denoted as
@@
where $(v,v^{\prime})\!\in\!\mathcal{E}\!\setminus\!\cup_{b\in\mathcal{N}^{\text{cav}}}P_{b}$ denotes an edge located on the GPLs. Note that $\Delta{N}^{\text{hv}}_{(v,v^{\prime})}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
Subsequently, by the BPR model in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right),$ (11)
where $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow computed by (9).
@@
The traffic flow monitoring windows are defined by $\Delta{T}$ in (4) for joint DLs and by $\Delta{\tilde{T}}$ in (8) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges.
@@ IV Dynamic Rerouting Approach
In this section, we propose a coordinated rerouting approach that dynamically optimizes route assignments for CAVs based on real-time traffic data. The objective is to ensure that buses adhere to their scheduled arrival times at each stop while enhancing the overall traffic efficiency of CAVs. The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and determining the optimal alternative routes for those CAVs.
@@ IV-A Identification of CAVs for Rerouting
where $\tau_{b,k-1}^{0}$ is the free-flow travel time from $v_{b,k-1}$ to $v_{b,k}$.
@@
$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right).$ (14)
Figure 3: Rerouting scheme at each intersection in DLs.
As illustrated in Fig. 3, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$. For any system time $t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]$, the set of CAVs on edge $(v_{b,k-1},v_{b,k})$ that need to be rerouted, denoted as $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, is expressed as
@@
As defined (6), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered. Namely,
$t_{\text{sys}}^{\prime}\in\left\{t\in[t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t)\geq(1\!+\!\lambda)\tau_{b,k}^{0}\right\}$ (16)
where $\lambda\!>\!0$ is a small tolerance.